The length of a rectangular playing field is 5 ft less than twice its width. If the perimeter of the playing field is 230 ft, find the length and width of the field.
Answer by checkley75(3666)

L=2W-5 IS THE LENGTH
W=WIDTH
2(2W-5)+2W=230
4W-10+2W=230
6W=230+10
6W=240
W=240/6
W=40 FT. IS THE WIDTH.
L=2*40-5=80-5=75 FOR THE LENGTH.
PROOF
2*75+2*40=230
150+80=230
230=230